OpenStudy (anonymous):

The temperature in degrees Fahrenheit can be expressed by the function F(c)=9/5c+32 where c is the temperature in degrees Celsius. Find the temperature in degrees Fahrenheit (to the nearest degree) if it is 31 °c outside.

OpenStudy (lalaland_lauren):

\[\frac{ 9 }{ 5(31) }+32\] \[\frac{ 9 }{ 155 } + 32\] 17.2+32 = 49.2